OpenStudy (anonymous):

Consider this reaction. 2SO2g+O2g yields to 2SO3g. What volume, in milliliters, is required to react with 0.640 g of SO2 gas at STP?

OpenStudy (anonymous):

We can use the following formulas \[m = n M\] where m is the mass, n is moles, and M is the molar mass, so figure out the molar mass of SO2,. And we also have to use another formula \[n = \frac{ v }{ V_{STP} }\] we can relate this with our first formula by doing solving for n first. \[m = nM \implies n = \frac{ m }{ M }\] now we can plug this into our second equation \[\frac{ m }{ M } = \frac{ v }{ V_{STP} } \implies v = \frac{ m V_{STP} }{ M }\] note that \[V_{STP} = 22.4 L/mol\]

OpenStudy (anonymous):

Thank you for taking time to answer

OpenStudy (anonymous):

Your welcome :)